Bobbi Reutter invested some money at 7% and 100 less than at 8%. Her total annual interest was $58. How much did she invest at each rate?
Answer by mananth(16946)

x= Amount invested at 7.00%
y= Amount invested at 8.00%
7%x+8%y=58
multiply by 100

7 x + 8 y = 5800 .............1

1 x -1 y = 100 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 8
7 x 8 y = 5800
8 x -8 y = 800
Add the two equations
15 x = 6600
/ 15
x = 440
plug value of x in (1)
7 x + 8 y = 5800
3080 + 8 y = 5800
8 y = 5800 -3080
8 y = 2720
y = 340
x= 440 Amount invested at 7%
y= 340 Amount invested at 8%
m.ananth@hotmail.ca